find the width of a rectangle with a length of 6 yd and a perimeter of 18 yd

Answer by richwmiller(17219)

what is the perimeter?
it is a walk around the edge of the rectangle.
the perimeter of a rectangle =2L+2W
we know the length (L) and we know the perimeter.
so we plug those in the formula
2L+2W=P
2(6)+2W=18
12+2W=18
subtract 12 from both sides
2W=6
divide both sides by 2
w=3 yds =width